### v1

# A blueprint for large-scale quantum-network deployments

Abstract
Quantum Communications is a field that promises advances in cryptography, quantum computing and clock synchronisation, among other potential applications. However, communication based on quantum phenomena requires an extreme level of isolation from external disturbances, making the transmission of quantum signals together with classical ones difficult. A range of techniques has been tested to introduce quantum communications in already deployed optical networks which also carry legacy traffic. This comes with challenges, not only at the physical layer but also at the operations and management layer. To achieve a broad acceptance among network operators, the joint management and operation of quantum and classical resources, compliance with standards, and quality and legal assurance need to be addressed.
This article presents a detailed account of solutions to the above issues, deployed and evaluated in the MadQCI (Madrid Quantum Communication Infrastructure) testbed. This network is designed to integrate quantum communications in the telecommunications ecosystem by installing quantum-key-distribution modules from multiple providers in production nodes of two different operators. The modules were connected through an optical-switched network with more than 130 km of deployed optical fibre. The tests were done in compliance with strict service level agreements that protected the legacy traffic of the pre-existing classical network. The goal was to achieve full quantum-classical compatibility at all levels, while limiting the modifications of optical transport and encryption and complying with as many standards as possible. This effort was intended to serve as a blueprint, which can be used as the foundation of large-scale quantum network deployments. To demonstrate the capabilities of MadQCI, end-to-end encryption services were deployed and a variety of use-cases were showcased.
Index Terms:
Quantum networks, optical networks, quantum key distribution, quantum communications.
I Introduction
There is a growing interest in the use of quantum-enhanced services, mainly based on quantum key distribution (QKD) [1, 2]. However, this requires the ability to manipulate signals at the quantum level and, consequently, the deployment of new technologies with very demanding technical requirements.
The principles of quantum mechanics imply that the measurement of a quantum signal alters it. Passive networks, without active elements such as amplifiers, must be used for its transmission. The propagation of signals at the single quantum level in a physical media—typically the optical fibre—entails an absorption probability that dramatically limits its range. If that medium is shared, other quantum or classical signals may limit the range even more. In the case of optical fibres, deployment is typically the most expensive part and co-propagation is essential. Until quantum repeaters are developed [3], it is necessary for QKD networks to handle these limitations.
Equally relevant to the success of quantum communications is the provision of the quantum-enhanced services. This involves the joint operation and management of the quantum and classical resources, which is an enabling technique of low-level ideas, like sharing the underlying physical infrastructure. Other important aspects for widespread deployment are the interoperability of deployed systems, standardisation, certification of quantum networks and devices, compliance with regulations, and quality assurance. MadQCI [4] is designed to be heterogeneous at the quantum as well as the classical network level, to comprehensively address these problems. It is also large and complex enough to significantly advance the state of the art. To date, with three years of continuous operation, MadQCI is the most complex, largest, and longest running testbed for quantum-communication technologies in Europe. To realise a diverse network and perform a comprehensive investigation, multiple technologies from various vendors were installed. The 26 QKD modules came from three different manufacturers, some using continuous-variable QKD (CV-QKD) and others discrete-variable QKD (DV-QKD). Table I summarises the main features of the network.
TABLE I: Summary of MadQCI features by domain.
Domain Nodes QKD links Total link length QKD modules Encryp.
TID 3 3 + border 38.4 km 7 3
RM 6 6 + border 91.4 km 9 2
Co-located 7 all-to-all 98.7 km 10 6
The complexity of the network was tackled using several administrative domains. The infrastructure was hosted in production premises of two network operators: Telefónica Innovación Digital (TID), the research and development division of the incumbent telecommunications operator in Spain, and the research and education network in the region of Madrid, REDIMadrid (RM). In addition, a third domain was used to handle the testing of a switched QKD network. This specific sub-network is referred to as the “co-located domain” and contains all the devices [5] that support switching. The network configuration resembles production networks—including the inter-domain border link—re-using already installed resources and adding new ones as required.
Regardless of the domain, the operation and management of the overall quantum network was based on a software-defined approach [4] to quantum networking. In the software-defined networking (SDN) paradigm, forwarding-plane resources are programmable and expose their capabilities through common and open interfaces [6]. In this way, the operation and management system not only configures and commands the resources, but also achieves the overall programming of the network for application-specific behaviour and performance. This design is very flexible and allows an easier integration of quantum resources in existing classical networks. In MadQCI, this paradigm enabled in a transparent way the operation and management of both the quantum and classical equipment required for quantum-secure communications and services.
As a result, cryptographic keys could be distributed between any pair of nodes in the network, independently of the QKD manufacturer and network operator. These capabilities were exposed to applications through standardised interfaces, which benefited from the programmability of the network as a whole. The outcome allowed the execution of a wide variety of use cases [4, 7, 8] fed with quantum-distributed keys. Specific hardware was installed to provide encryption at OSI levels 1 and 2. Software applications were used for test on higher OSI layers, such as an IPsec implementation [9] and one based on TLS [10].
The paper is organised as follows. In the next sections, the related work and an overview about quantum and optical communications is outlined and the coexistence and integration approach of classical and quantum optical networks is discussed. Then, a detailed description of the solutions implemented in MadQCI and its main performance are given to serve as a blueprint for large-scale quantum network deployments. The paper finishes with conclusions and an outlook.
II Related work
Quantum cryptography has been the first major technological expression of quantum communication and QKD grew around the idea of using optical communications and networks since its earliest days. Examples include the Boston testbed [11], which already involved optical networking techniques such as switching or dense wavelength division multiplexing (DWDM). Other well-known examples are the Vienna [12] and Tokyo [13] quantum networks, which used dark optical fibres deployed in the field. The first iterations of MadQCI also date back to that time [14], along with others such as the Hefei testbed [15].
Most of these practical exercises focused on showing the capabilities of quantum-communication developments, rather than on how to install these into existing infrastructures [16, 17, 18, 19, 20]. In contrat, the efforts in MadQCI were aligned with this last goal since its early days. Novel optical networking approaches to quantum communication were proposed [21] and it was shown how a pioneering commitment to a software-defined approach to quantum networking enables a fast and flexible convergence of quantum and classical communication in production environments [22, 23]. Furthermore, the MadQCI team is a leading actor in the standardisation of QKD communication at ETSI ISG QKD [24], which promotes a dialogue between industrial actors such as manufacturers and network operators.
In parallel, numerous technological improvements over the last decade have brought QKD to commercial status. Deployments such as the 2,000 km Beijing-Shanghai QKD link [25] and the 4,600 km long QKD space segment [26] have demonstrated the feasibility of quantum secured broad backbone networks. Other examples demonstrate the penetration of these technologies in specific niches, such as the ability of the Hefei quantum network to efficiently encrypt its institutional telephone traffic [27] or the efforts in Bristol to securely transport subscriber databases in 5G [28].
Since 2019, a broad effort is underway to establish he European Quantum Communications Infrastructure (EuroQCI), a pan-European quantum network based on the results of several projects funded by Quantum Flagship [29] and Digital Europe Programme [30] research programmes, among others. Examples of these projects are OpenQKD [7], CIVIQ [31], and the current QSNP [32], which strongly contribute to the discussions in this paper. Currently, there are several efforts to deploy quantum communication infrastructures (QCI) in Europe. A list of these can be found in [33].
The next big step in the adoption of quantum communication is the leap to the space segment, since it would enable worldwide quantum connectivity. Also, further progress is being made in other types of quantum communications such as quantum entanglement distribution—to enable the so called “quantum internet” [34]—or other quantum-enhanced applications such as quantum clock synchronisation [35]. Communication infrastructures are being prepared to support any kind of quantum communication, beyond quantum cryptography.
III Quantum and optical communications overview
This section introduces some topics in telecommunications and quantum technologies, particularly those transmitting in optical fibres.
III-A Production networks
When considering the embodiment of quantum technologies in optical networks, it is necessary to build on a suitable model, which may differ depending on the nature of the requirements [36].
The current networking landscape relies on several heterogeneous communication systems with different design, technology, and service fundamentals. Besides, each of them has a different degree of interoperability, often limited by the commercial strategy of the vendors. Examples are the more traditional Ethernet-based local area networks; the passive optical networks (PONs) for home access; or the optical networks for aggregation and backhaul, as Fig. 1 depicts. All operate over optical fibres and can carry the same type of traffic. However, they have different requirements for the deployment of new systems. In addition, network operators and service providers must comply with legal and regulatory requirements—e.g. critical infrastructure protection, lawful interception, and secrecy of communications—that prevent the deployment of novel devices without certification and industry-wide standards. The deployment of quantum communication adds additional requirements to these already overloaded designs, which have to be integrated to deliver cutting-edge services.
Figure 1: This figure illustrates the complexity and diversity of optical communications. On the left, two access networks are shown, an enterprise and a home network based on passive optics. In the centre, two (highly simplified) nodes of a typical optical network are depicted, one dedicated to the aggregation of multi-purpose subscriber traffic—and possibly with FOADM optics—and one backbone for handling large volumes of traffic, typically with equipment such as ROADMs. On the right is a data centre of a digital service provider, operating over-the-top of the network operator. In terms of subscriber provision, the optical network is delivering two services: [1] provision of lightpaths and [2] carrying and processing of traffic. The target is to integrate quantum communication technology with all types of traditional (classical) optical fibre-based communication.
The different industrial and commercial approaches should also be taken into account when modelling the optical networks. Network operators focus their activities on the transport of information between their subscribers. Thus, their network design must deal with the complexity of equipment spread across customer sites and central offices, while assuring the aggregation and transport of multi-service traffic. Whereas, in the case of digital service providers, the complexity lies in their multi-tenant, distributed data centres, which are interconnected “over-the-top” of the network operator, abstracting the complexity of many parts of the network. The conception of quantum communications and the service they provide can be fundamentally different in these scenarios, and involves variables ranging from regulatory issues to how or to whom to bill for these services.
A description of a network can be achieved by explaining its network functionalities, usually classified in functional planes—i.e., the forwarding, control, management, application, etc. planes—. The previously explained context bounds how, when, and where those functionalities are executed. In this regard, a production network cannot be modeled as a naive graph of connected nodes. It is common for forwarding-plane functionalities, i.e., those functions that enable the transport of information over the different transmission media, to reside at all nodes of a network. However, the majority of the control-plane functionalities may reside in a single centralised component. In this case, the unavailability of a single node may compromise the connectivity of the entire graph. Furthermore, other elements, such as the facilities that physically host the forwarding devices and provide security and privacy to the system, are also functional to the network. Therefore, a more appropriate model should include which network functionalities are performed on which element, with a focus on the specific end use or application.
This paper attempts to incorporate these insights into discussions on the coexistence of classical and quantum technologies.
III-B Optical networks
From a technological point of view, an optical network is a communication infrastructure capable of providing lightpaths. In this sense, circuits are more important than packets or datagrams for this model. Each lightpath carries optical signals that encode subscriber traffic. In some cases, a lightpath may be managed and operated directly by the customer—e.g., a dark optical fibre is leased by a digital service provider to connect some remote data centres—.
The optical network typically uses wavelength division multiplexing (WDM). Although a bidirectional approach is possible, duplexing is usually achieved by using pairs of optical fibres, one for receiving and one for transmitting. Transceivers encode the subscriber traffic in optical signals for specific duplexing and WDM channels.
Multiple elements such as filters, multiplexers—e.g., optical add-drop multiplexers (OADM)—, cross-connection systems, etc., enable the steering of these signals. In order to achieve long range transmissions, signal re-amplification, re-shaping or re-timing techniques are implemented. In many cases it is not possible to separate these elements from the frame they carry. For example, a muxponder that aggregates and de-aggregates several information streams to optimise the transmission of large volumes of data may only work for a particular technology—for example, those specific to SDH/SONET—.
Finally, several visions or generations coexist in the industry to achieve a higher or lower reconfigurability of lightpaths, as also depicted in Fig. 1. A common approach is called “FOADM”, which stands for fixed OADM. In these designs, elements such as fixed filters and multiplexers prevail, i.e., without the possibility of changing their frequency selectivity. Re-tunable transceivers may be used, depending on the connectivity requirements in the assembly. The “ROADM” approach uses precisely reconfigurable OADMs or analogous elements. This means that the optical mesh can be reconfigured by means of frequency selective elements to establish new lightpaths on demand. The degree of meshing may not be all-with-all, and hybrid approaches are possible—e.g., FOADM in the access network and ROADM in the backbone, or vice versa—.
The result will be a meshed network capable of transmitting and receiving both information flows and lightpaths between nodes. The degree of meshing, the degree of reconfigurability of the lightpaths, the use of specific optics—such as passive optics for home access—, among other decisions, is determined by the design. The preferred design depends on the trade-off between performance—which includes not only throughput, but also latency, connectivity, availability—and cost.
III-C Quantum key distribution and cryptographic applications
The main quantum communication technique currently exploited is QKD, due to its level of maturity. However, experiments are also being conducted with other primitives of quantum cryptography, e.g., quantum oblivious key distribution for performing oblivious transfer [8].
QKD provides synchronised sources of symmetric cryptographic keys between two remote systems [1, 2]. This technique relies on quantum physics to deliver information-theoretically secure (ITS) keys to a final cryptographic application [37], as Fig. 2 shows. Indeed, an ITS cryptosystem cannot be broken even if the adversary has unlimited quantum computing power together with the best possible algorithms. Therefore, in theory, QKD generates highly secure cryptographic keys, immune to any algorithmic cryptanalysis, even those that may arise in quantum computing and which do threaten classical key-exchange systems [38].
Figure 2: Main scheme to deliver end-to-end quantum-secured cryptography using QKD. The symmetric quantum-distributed key is delivered to both ends of a cryptography system. For this, either a point-to-point QKD link exists, or an approach such as having trusted nodes that allow hop-by-hop key forwarding is used.
Since optical quantum channels suffer from a series of limitations regarding throughput and range, a network is needed to provide any significant end-to-end service. The quantum information is carried embedded in quanta of energy which, by their nature, are extremely sensitive to interactions. A widely used approach to achieve end-to-end key distribution is using “trusted nodes”, intermediate nodal sites where forwarding operations can be carried out without compromising the security. In each node, a component named “forwarding module” (FM) performs a key transport protocol, consuming the generated key and using another ITS primitive—e.g., Vernam encryption [39]—for enciphering and deciphering the final distributed key in each hop.
This enables end-to-end quantum-enhanced services both between remote subscribers and optical network itself—e.g., encryptors [40] for protecting the subscribing traffic—. In practice, cryptosystems based on QKD are not ITS. Any implementation requires assumptions which may not exactly reproduce a theoretical model. All elements of the network as a whole must be as secure as possible.
III-D Management and operation. Software-defined networking
Central to the evolution of telecommunication networks is their operation, administration, and management functionality, as it determines the overall performance of the systems [36]. They are also essential to achieve the interoperability between systems.
As described above, current networks are heterogeneous, based on different technologies, generations or vendors, but their management should be integrated. Traditionally, the network management and operation is achieved using techniques such as signalling channels or the direct configuration of devices—e.g., with SNMP [41]—.
Software-defined networking enables network programmability for enhancing the service provided [6, 42]. Traditionally, network equipment was design to save costs by limiting its configurability and focusing on performing specific tasks. Modern approaches promote the supply of more open and general-purpose equipment, often making use of software engineering techniques, such as exposing its functionality through software interfaces or domain-specific programming languages. This allows decoupling of functional planes, both the control and forwarding planes as well as the application plane; component disaggregation; and other software-enabled techniques such as virtualisation. In most cases, the SDN design assumes a logical centralisation of the control functions into a specific component called controller, which manages the underlying general-purpose infrastructure. The result is a network that can be programmed to meet the requirements of each application, either in concrete sense—e.g., each network usage per subscriber—or abstract—high-performance usage or low-energy impact usage—.
Analogously, quantum networks can adopt this approach. In fact, this work discusses and applies the concept of the “quantum forwarding plane” (QFP): functionalities that enables the end-to-end quantum-enhanced provision. This includes the FMs that enable key forwarding in trusted nodes, as well as quantum information systems—e.g., QKD systems—and other infrastructure elements. All the elements in the QFP may be controlled by an unique control plane. This will be fundamental for the integration of quantum technologies. A proper and joint management of classical and quantum resources can address low-level issues, such as sharing the underlying infrastructure, as well as other issues already mentioned, such as security, legal requirements, and quality assurance.
These and other issues are the subject of standardization, which is a fundamental task for QKD networks. The work described in this manuscript benefits from the standardisation activities on SDN in quantum networking [22, 23] driven by the ETSI GS QKD industry specification group.
In this regard, the ETSI GS QKD 015 specification [43] defines a software-defined node in a QKD networks as a set of components, including two forwarding elements: the QKD modules, which handle quantum signals and protocols, and the set of local key management systems (LKMS), that securely manages and forwards the key. Fig. 3 depicts this architecture. This specification also determines an information model that enables controlling every nodal component in a software-defined way. The subscribers may request end-to-end QoS-driven QKD services by ETSI GS QKD 004 [44], passing the use-specific requirements to be fulfilled by the network. Other specifications, such as ETSI GS QKD 018 [45], enable inter-domain orchestration or key management inter-operation. The architecture of the network elements is also being standardized to include the definitions of the QFP and the FM used above [46].
Figure 3: The architecture of a node in a software-defined QKD network as specified in the ETSI GS QKD 015 specification and implemented in MadQCI.
Optical networks can be operated and managed this way using so-called SDN protocols—e.g., NETCONF [47]—, YANG modelling language [48], and information models such as Transport API [49] or OpenROADM [50]. This approach enables the necessary compatibility to operate and manage both classical and quantum resources in a joint manner.
III-E Optical quantum networks
This section connects what was described so far to discuss the coexistence and integration of quantum communications in optical networks.
Quantum signals are of very low power and sensitive to interactions [17]. When transmitted in a shared optical fibre, the Raman effect and four-wave mixing of adjacent channels can cause disturbances in the quantum channels [51, 52]. In consequence, the management of the quantum and classical channel allocation is fundamental [53]. Active components, such as signal amplifiers, are destructive for the quantum nature of signals, regardless of the re-amplification, re-shaping or re-timing technique that they implement. The possibility of using other signal reconditions of passive nature, such as polarisation or dispersion compensation, may depend on the nature of the quantum communication.
In optical networking, multiplexers and other elements are not perfect and its effects, e.g., crosstalk between channels, signal reflections, and backscattering might influence the guardband communications [17, 54, 55]. As a result, coexistence must be implemented with optical equipment with good inter-channel isolation and limited launch power of the classical signals. Networks adopting the ROADM approach often use equipment that integrates amplification by design, such as wavelength selective switches (WSS). The quantum channels have to be dropped, filtered and added outside of the WSS to comply with the requirements. These restrictions have a direct impact on both the design and technology selection for the network.
When adopting such solutions, the requirements of the classical signals need to be managed for protecting the quantum channel. The power control negatively influence the propagation of classical signals. In order not to disturb a quantum system, it might be necessary to reduce the launch power of a classical channel so that the received power falls below the sensitivity of the classical receiver. The requirements could be met with better equipment or additional amplification on the receiver side once the channel has been dropped.
For many of these problems, as noted above, interoperability between the control and management functionalities of all elements becomes critical. If the operation and management system for the quantum resources can interoperate with that of the optical infrastructure, information can be consolidated and decisions can be made to increase performance. However, it is fundamental that industry-wide standards are developed to facilitate such coexistence and integration.
III-F Coexistence and integration approach
To conclude the discussion, a successful approach to integrate quantum elements in optical networks is presented, which is the one followed in the deployment of MadQCI. Two abstractions have been used primarily: functional planes, which are very common in SDN discussions, and network components, used in architectures such as the one described in ETSI GS QKD 015. Other abstractions are possible, such as describing the QKD network in terms of its layers—e.g., quantum, key management and communication layers [13]—, but this work does not do so.
In this regard, the ETSI GS QKD 015 architecture focuses on the control plane and abstracts other elements. It explicitly includes the QKD modules and the controller-agent assembly, but it does not include other elements that enable them, such as the shared optical network or the IT systems. Furthermore, that specification does not deal explicitly with the FM. Therefore, this work provides a more detailed architecture than the ETSI GS QKD 015 specification.
The design of each node will be the set of components that performs all the network functionalities, both classical and quantum. In quantum cryptography, this involves both the key management functionalities and those that enable the networking. An example of the former can be a hardware security module (HSM) for the secure handling of cryptographic keys or the management of different subscribers. This type of equipment exists in other fields and is subject to highly demanding standards, so security is not compromised [56]. The latter include the QFP components that enable a controlled and managed forwarding of both quantum signals and quantum-distributed resources. Fig. 4 shows a general scheme of the architecture presented in this work, based on a set of components that allows quantum and optical to coexist.
Figure 4: The image shows the set of components that would enable the coexistence of quantum and classical communication technologies in optical networks. It includes all the elements necessary for the provision of the service. It is relevant that the forwarding module (FM) that transports the key has been represented as an infrastructure component. Despite routing not being a key management functionality [56], it has been usually assumed to be part of the LKMS in the QKD literature. Note that the switch depicted at the bottom is a fiber optic switch. A set of them was installed in the co-located domain to enhance the switching capabilities of the optical network.
From top to bottom, the assembly is arranged to enable both hardware and software-based subscriber traffic, and all of them will require quantum-distributed keys to the network. This distribution is made possible by components that control and manage both, the key and infrastructure components of the network. In a complex case, this will be done in an integral way, with a holistic view of the network. This infrastructure can be a heterogeneous set of quantum systems, together with the FMs, as well as the optical network shared by all the other components. With programmatic interfaces, such as those defined by ETSI GS QKD 004 or ETSI GS QKD 018, applications can request enhanced services from the network thanks to an heterogeneous set of resources.
The result is a node ready to integrate an optical network with quantum technologies focusing on cryptographic applications, mainly based on QKD.
IV MadQCI, a blueprint for quantum networking
This section describes MadQCI to serve as a model for other QCI deployments. The fundamentals described so far were taken into account during its design and deployment, so it is a particularisation of those. The aim is to detail the problems that were addressed and the solutions applied to provide a blueprint for any large-scale quantum network deployment.
In particular, this detail concerns the three years in which the testbed was running during the OpenQKD and CIVIQ projects. Its scope and complexity make it possible to provide a detail that had not been achieved so far in the state of the art, as it was the most complex, largest and longest running testbed for quantum communication technologies in Europe. Fig. 5 depicts the layout of MadQCI during the period detailed.
Figure 5: MadQCI overview. 26 QKD modules from different manufacturers were installed in the 9 production facilities of RM and TID. More than 130 km of fibre optic pairs supported both classical and quantum signals including, in some cases, third party traffic. Three OSI level 1 encrypted links were deployed, as well as 6 level 2 Ethernet encrypting devices. A single pair of optical fibres was used in each inter-node connection, but different schemes of coexistence were implemented. The link between Quevedo and Norte was co-managed as a border connection.
Additionally, the key and network operation and management was performed as much as possible in an SDN-aware manner. The approach followed was based on the general design outlined in section III-F, but devoted to SDN techniques and technologies. The level of integration with each component depended on its level of SDN-awareness.
The network supported several use cases and demonstrators of different projects. Specific hardware was installed to provide encryption at OSI levels 1 and 2 (ADVA and R&S respectively). A level 3 IPsec [9] and TLS [10] software implementations were also tested. The result allowed the execution of a wide variety of use cases [4, 7, 8] enhanced with quantum-distributed keys.
IV-A Premises and equipment
The uniqueness of MadQCI is that it resulted in a heterogeneous network installed in real-world production optical networks, with multi-vendor equipment and meeting the requirements of the two network operators involved.
IV-A1 Classical-specific equipment
REDIMadrid is the research and education network in the region of Madrid and provides connectivity services to several Madrilenian universities and research centres. Its infrastructure is meant for delivering IP-based connectivity services and it is outfitted with production-grade optical network devices—e.g., Ciena 6500 or Adva FSP 3000—. RM facilities are mainly central offices or communication rooms, which fulfilled the security requirements of trusted zones. The network implements path diversity and redundant channels, so that quantum integration could be trialed without risking the service delivery.
(a) ID Quantique QKD modules, one receiver and one transmitter for performing the key forwarding. One operates at 1310 nm and the other at 1550 nm (ch. 34). A Ciena WDM multiplexer can be seen on top.
(b) Huawei QKD modules, one receiver and one transmitter. Each module consists of a two height units optical box and a one height unit server for the key distillation. The set of Huawei experimental devices were deployed along the two TID and RM domains, as a sub-network, the co-located domain. All network equipment was shared. R&S Ethernet encryptors were deployed at each node of this co-located domain except Quirón.
(c) An ADVA FSP 3000 optical network device (below) and a BiDi ADVA WDM shared equipment (above) that enabled traffic coexistence. The ADVA device was outfitted with 3 insertable modules (right half); the central one supports RM production traffic while the lower one (with yellow single-mode fibres connected) was a 5TCE-PCTN-10GU+AES10G module for ciphering the traffic using the quantum-distributed key. Then, the traffic of the modules was multiplexed in the BiDi ADVA multiplexer.
Figure 6: Different examples of the equipment deployed in MadQCI, in the Quijote node in this case. All devices displayed here were mounted in the same 19” rack cabinet.
The premises available in the Telefónica Innovación Digital domain were production central offices. These facilities enabled the daily operation of communication services by connecting the wired and radio subscriber access networks to the backbone transport networks. As such, they were much more regulated critical infrastructures—e.g., taking pictures of the equipment once installed was not authorised—, but suitable for validating quantum-safe communication solutions. This is also why the MadQCI network was designed in this segment from scratch and used dedicated equipment and non-illuminated fibres.
In both network domains, the quantum-communication technologies, both quantum and networking, were installed in regular 19” rack cabinets and had a power supply of 240 V, which fits the industrial standards. At the Norte and Concepción nodes—i.e., central offices of the TID domain—power inverters were used to adapt the direct current typically used in these facilities. In all the premises, there were regular refrigeration conditions. The optical transmission media were standard fibres—e.g., ITU-T G.652—and production-grade connectors were used, such as LC or SC with different polishes. Fig. 6 shows several examples of the deployed quantum and classical equipment.
For delivering the quantum-safe services, both software and hardware resources were deployed. The hardware devices were network encryption equipment consuming the quantum-distributed to cipher the subscriber traffic while performing their network tasks. High-security Rohde & Schwarz SITLine ETH encryptors were deployed for supporting ciphered L2 connectivity services. Analogously, some ADVA 5TCE-PCTN-10GU+AES10G modules were installed to enable L1 encryption [40]. Regarding the software applications, the domain-specific use cases were demonstrated using the same IT systems that ran the SDN-like control and management system, described in the section IV-C.
IV-A2 Quantum-specific equipment
Three different QKD vendors were present in MadQCI, namely Huawei, ID Quantique and Toshiba. An additional, experimental system, from AIT was also installed remotely in the network as a further demonstration of the integration capabilities of the network. The ID Quantique and Toshiba discrete-variable QKD systems were deployed in both domains. The former ones host in the same chassis several blade-like subsystems, such as the quantum devices itself or optical equipment for the channel adding and dropping. The latter ones were designed to be connected in the middle of a fibre pair that connects a facility, in the nearest point to the ducts. It essentially by-passes every incoming or outgoing optical signal, with some losses, except for its own quantum and signaling channels. The co-located set of Huawei continuous-variable QKD systems rely on external optical equipment, shared with the ID Quantique systems.
It should be noted that QKD technology is currently establishing itself in the market with multiple business models, which conditions their integration. Many manufacturers choose point-to-point solutions, such as HEQA, ID Quantique, QTI, ThinkQuantum or Toshiba in DV-QKD or KEEQuant and LuxQuanta in CV-QKD. Each one shows its specificities; for example, the way the Toshiba systems multiplex channels or the HEQA equipment that integrates a receiver and a transmitter to facilitate the key forwarding. Other business models take advantage of the new QKD techniques to offer other competitive advantages, such as Q*Bird, which uses systems based on measurement-device-independent QKD to offer a configuration based on a central hub with multiple emitters arranged in a star configuration. A potential QKD breakthrough will come when miniaturisation in on-chip photonics enables other form factors, such as small form-factor pluggable (SFP) integrations.
Regarding quantum random number generation systems, the integration of Quside systems was tested remotely at MadQCI. In any case, there are also resources from other manufacturers, such as ID Quantique or Quantum-Industries, with different technical approaches. For quantum networking, there are two main ones; those based on PCI cards for IT systems and rack-mounted solutions. In both cases, the delivered entropy rates are similar.
IV-B Quantum and classical coexistence
This section details the deployments that enable quantum connectivity in each of the links. Note that it will be these QKD links that will enable the provision of the SDN-like end-to-end service that will be explained in the following section IV-C.
For achieving the quantum and optical integration, both quantum and classical channels were allocated in the DWDM grid using the most suitable equipment in each case as stated so far. However, in MadQCI several coexistence schemes were tested, both for quantum-classical and quantum-quantum coexistence. Fig. 7 shows these main coexistence schemes implemented in MadQCI.
Figure 7: Main coexistence schemes used in MadQCI. a) Duplex usage of a fibre pair with QKD and classical signals multiplexed in both directions. Several multiplexing stages are used to manage losses, but a single one could be used. The excess loss is compensated with an amplifier at the classical signal input. b) A bidirectional scheme is implemented with a BiDi-multiplexed fibre devoted to classical channels and the other to quantum signals. An additional approach, not depicted here, is using circulators and an unique fibre strand for a bidirectional coexistence [57]. C) Example of a switched QKD network as implemented by Huawei in the co-located domain. On the left, ROADM-type switching is achieved by combining optical fibre switching with multiplexing (enclosed by a dashed line). On the right, FOADM-style switching, where it would be sufficient to transmit on one or the other channel to reach different destinations. Finally, the classic part shows how the wide receive dynamic range of optical transceivers is exploited to pair a couple of transceivers with different channels, which allows better handling of guardbands.
In the RM domain, the legacy production equipment could be used for aggregation and routing of signals to the terminal equipment. Depending on the case, it was necessary to install ad hoc equipment for the final stages—e.g., a demultiplexer for splitting two quantum coexisting channels—. In addition, some QKD systems operated at 1310 nm instead of the 1550 nm window. Regarding the active components, any classical signal recondition was performed after the signal aggregation, e.g., placing the amplifiers near the transceivers. RM operates an IP-based network and the IT systems that supported the use cases were provided with connectivity by RM network.
In the TID domain, the integration could be more flexible, as the equipment was not shared with other transmissions. In these three nodes a higher density of equipment was deployed and coexistence was more intensively tested. Attack or eavesdropping emulators were additionally installed. Several connectivity technologies were also tested: switch-to-switch Ethernet connectivity using colored transceivers, both in the clear and encrypted (R&S encryptors); OTN-like muxponders by Huawei (OTU4 graming); and the multi-frame optical encryption by ADVA modules. On top of these, network logic was configured to ensure visibility between all QKD, IT and encrypting systems using L2/L3 network devices—e.g., ADVA FSP 150-XG304 or Huawei Quidway S6700 Series—.
This network connectivity is fundamental in QKD since, together with the quantum resources allocated in the optical network, the cryptographic key material have to be securely managed and transported, as described in section IV-C. Thus, there are different components that must co-operate for delivering the quantum-enabled service. However, this co-operation must happen in the trusted scenario required to ensure the information-theoretically security. Secure design principles such as economy of mechanism or access control techniques, as L2/L3 slicing, helped to achieve this.
Additionally, a border link was implemented between both domains to enable testing cross-domain use cases. An ID Quantique QKD system was deployed there and IP connectivity between the IT systems of both domains was provided in the last stage of the iteration. Please, note that the co-located domain, namely the Huawei and R&S devices, was deployed in both networks since the very beginning.
Regarding the co-located domain, Huawei implemented a switched QKD network [5]. The objective was to use transmitters and receivers between any node in the network as needed, and with a lower number of modules installed. In addition, this approach makes it possible to avoid the need for trusted nodes in low-loss scenarios, such as metropolitan or access networks with limited reach. For this purpose, an approach closer to the ROADM style was followed.
Each node was equipped with a fibre optic switch between the fixed optical multiplexers which, in this case, allowed channels 34, 37 and 38 to pass through. This performance of the fixed multiplexers combines with that of the fibre optic switches to achieve a cross-connection comparable to that of a network based on ROADMs but without the amplification typical of the WSSs. In addition, CV-QKD systems could tune their lasers to any DWDM channel, so any connection could be established if a sufficiently low-loss optical path was available and, at the nodes in the path, any of the three channels were not being used by another connection.
As a result, this co-located domain was able to test a total of 34 possible routing combinations between the 10 CV-QKD modules installed on 7 nodes.
Table II aggregates the connectivity achieved on each of the links. Thus, they will be available to the operation and management system described in the next section to provide the end-to-end quantum-enhanced services.
TABLE II: Connectivity achieved on each of the network links, so that they can be exploited by the SDN-like operation and management system as explained in the IV-C section. All domains use the same fibres (and usually the same optical equipment). Distillation and synchronisation channels of the QKD systems are included as classical channels, as well as encrypted channels (e.g., in the Quintín-Quijote link there are 6 classical channels, 2 carrying encrypted information, and 2 quantum channels). The co-located domain (Huawei CV-QKD systems) is characterised using the same links although its all-to-all switching capabilities makes the overall results not comparable. In all cases, the contribution of the co-located domain to the amount of quantum channels is a single quantum channel per link. However, due to the switching capabilities implemented by Huawei, it is possible to find up to four CV-QKD channels in a single fibre [5]. The link losses of this co-located domain have been measured by the QKD systems and they include the intra-node losses.
Link Length [km] Loss [dB] Domain Vendor Keyrate [kb/s] Channels Details
Quintín-Quijote 24.1 7.47 RM ID Quantique 1.95 Class.: 6, Quan.: 2, Encryp.: 2
11.8 Co-loc. Huawei 6.2
Quijote-Quirón 24.5 8.72 RM Class.: 4, Quan.: 1, Encryp.: 1 Only deployed at
12.1 Co-loc. Huawei 3.3 co-located domain
Quijote-Quevedo 7.8 4.00 RM ID Quantique 2.04 Class.: 5, Quan.: 2, Encryp.: 1
4.8 Co-loc. Huawei 11
Quevedo-Quijano 33.1 11.80 RM Toshiba 37.21 Class.: 5, Quan.: 1, Encryp.: 0
Quijano-Quinto 1.9 2.0 RM Toshiba 2857.12 Class.: 5, Quan.: 1, Encryp.: 0 5G core simulator
Distrito-Norte 15.0 8.30 TID ID Quantique 1.47 Class.: 3, Quan.: 2, Encryp.: 1
10.2 Co-loc. Huawei 8.4
Distrito-Concepción 15.3 14.3 TID Toshiba 242.33 Class.: 4, Quan.: 2, Encryp.: 1
20.1 Co-loc. Huawei 0.11 via Norte
Concepción-Norte 8.1 2.9 TID Toshiba 1039.98 Class.: 6, Quan.: 2, Encryp.: 2
11.1 Co-loc. Huawei 7.4
Quevedo-Norte 3.9 3.0 RM, TID ID Quantique 1.38 Class.: 4, Quan.: 2, Encryp.: 1 Border link
4.3 Co-loc. Huawei 12
IV-C Management, administration and operation
In MadQCI, several management, administration and operation (OAM) mechanisms coexisted and were tested. Relevant to this paper is the software-defined management and operation based on ETSI GS QKD standards implemented at UPM, named SD-QKD Stack in this work. It is based on the approach explained in section III-D and depicted in Fig. 3. However, there were also some QKD vendor-specific key and network management systems deployed, as well as the management systems of the classical resources.
The SD-QKD Stack enabled the end-to-end quantum services by consuming the QKD keys generated by the links described in the previous section. It is composed of the SDN controller, the set of its agents and another set of local key management systems (LKMS) to perform the key management and forwarding tasks, according to the ETSI GS QKD 015 specification. This system therefore implements both key management and networking functionalities. Thus, a joint management of both network and cryptographic key resources was achieved, and a QoS-driven service as specified in ETSI GS QKD 004 could be provided [4].
It operates as follows. First, when a service is requested via the ETSI GS QKD 004 interface at two endpoints—or using the ETSI GS QKD 014 ones—, the agents at each node collected the necessary information and relayed it to the controller, that decided which settings are suitable in each case. Then, it commands the agents to configure the resources of each node, both the LKMS and the QKD systems, to provide the service. The needed keys generated by the QKD systems on each link of the calculated path are then collected, independently of their manufacturer, and managed by each LKMS. As this component includes the FM of each node, it consumes the keys to perform the key forwarding along all the required hops in a vendor agnostic way.
Regarding the classical resources, in general, the optical and connectivity infrastructure was manually or automatically managed and operated using the regular techniques in each domain. This included methods such as configuring the devices via command-line interface, using vendor-specific controllers or setting the configuration remotely.
In an ideal approach, this operation and management is joint with that of the QKD resources. Specifically, the ETSI GS QKD 015 specification defines some optical-networking related fields in its data model, thereby enabling this integrated software-defined control. For instance, the SDN controller is able to configure the set of SDN agents the required wavelength and expected attenuation in each hop, to configure these requirements in the SDN-aware devices, both optical and QKD. This shows the network programmability that SDN promotes, since those optical-networking requirements may arise from the quality of service requested by a subscriber using the ETSI GS QKD 004 interface, together with all other connectivity and key management requirements.
However, each system has a different degree of maturity, so such integration is an ongoing task and only specific experiments have been performed, as the following section V explains. Another approach is using other interoperability mechanisms, such as the optical-quantum orchestration promoted by the ETSI-GS-QKD-018 specification, that it is also an ongoing task.
V Performance
This section details some of the most relevant results, especially with regard to coexistence and integration with optical networks in production.
(a) Coexistence scheme that illuminates the Quijote-Quevedo link. It is based on the aggregation of the quantum and classical channels by means of two multiplexers in cascade, which minimises the optical losses of the former. The legend is the same as the one used in Fig. 7
(b) QBER and keyrate of the ID Quantique QKD system installed in Quijote-Quevedo.
(c) Spectrum of one line outgoing from Quijote to Quevedo.
(d) Spectrum of the same line ingoing to Quijote from Quevedo.
Figure 8: Coexistence in the Quijote-Quevedo link, outfitted with a Huawei CV-QKD link and a ID Quantique DV-QKD link.
V-1 General overview
In general terms, the installation and connection tasks were successful in every domain and had far exceeded the objectives of the associated research projects. Table II gathered the MadQCI main features and capabilities.
All designs were successfully deployed and only a few minor issues related to configurations or malfunctions of different equipment had to be resolved. Regarding the configuration, the usual complexity of any network deployment was faced, such as managing the network addresses of each device or configuring the keys or certificates for TLS-based connections. The most usual malfunction issues were related to the information systems attached to the network and QKD equipment, both those hosting the SDN software and those that run the key management systems of each vendor: permission and log management, software maintenance, hardware malfunctions… There was also a complete migration of a node to another installation—Almagro to Distrito site—, precisely because it was a production central office.
V-2 Quantum and classical coexistence
Regarding the co-propagation of quantum and classical signals, two examples are highlighted in this work. The most complex nodal schemes can be found in the Appendix VI.
The first one is shown in Fig. 8 and represents the performance of the link between Quijote and Quevedo nodes, in which two QKD links coexisted with classical data channels, both encrypted and in the clear—only the equipment supporting this link is showed—. The Huawei continuous-variable QKD link operated in the C band—channel 37 or 38 depending on the established link— while the ID Quantique discrete-variable QKD link was in the O band. The quantum-distributed key in both cases was delivered to the the SD-QKD Stack, to forward other key material and support the consuming applications. In this node, an Ethernet R&S encrypting device was installed, which is an example of these applications.
The other highlighted example includes subscribing channels. It is depicted in Fig. 9 and shows the performance of the Quintín-Quijote link. Again, only the devices which transmitted or received signals from that specific link are depicted. However, due to the switching capabilities of the co-located CV-QKD devices, it could exist quantum signals bypassing through the node. This link used a multiplexer for aggregating the signaling from two QKD links, Huawei and ID Quantique, both operating in C band—channels 34 for the DV link and 37 or 38 for the CV link—. In other MadQCI links, the Toshiba discrete-variable QKD systems operated at C band while the ID Quantique ones at both O and C bands—channel 32 the one not depicted—. For multiplexing the classical channels, a bidirectional multiplexed was installed, and supported the signaling channels of the QKD systems, the ciphered channels of R&S and ADVA encryptors and other RM third-party channels.
(a) Coexistence scheme that illuminates the Quijote-Quintín link. It is based on using one strand for classical signals, with a bidirectional multiplier, and the other one for quantum signals. The legend is the same as the one used in Fig. 7
(b) Performance of the L1 encryption by ADVA measured with Iperf tool. In theory, the cipher fills the frames with exactly 10 GbE of information for security, but this is not reflected in the host-to-host measurements due to the peculiarities of the connection, such as headers or TCP retransmissions.
(c) Spectrum of one line outgoing from Quijote to Quintín. In dashed line, the position of the quantum channels at the other fibre of the BiDi multiplexer for illustrative purposes (in this link there is no quantum and classical coexistence).
(d) Spectrum of the same line ingoing to Quijote from Quintín.
Figure 9: Coexistence in the Quijote-Quintín link, outfitted with Huawei CV-QKD and a ID Quantique DV-QKD links, both at C band. The spectrum shows the fibre dedicated to the classical data channels, both encrypted (L1 by ADVA and L2 by R&S) and in the clear. Some of the latest are RM third-party channels.
These are two examples of how the network was capable of delivering quantum-distributed keys in any link of the network while supporting third party communications in some links. For supporting QKD, both quantum and classical channels co-propagated over the network, including distillation and synchronisation channels and data channels for the uses cases. The SD-QKD Stack enabled the end-to-end connectivity, as showed in the following section.
V-3 Management, administration and operation
The operation and management of the quantum network resources and cryptographic key was performed by the SD-QKD Stack based on ETSI GS QKD standards described in section IV-C. The combined operation of the controller and the sets of agents and LKMS allowed the available resources to be operated in a holistic and end-to-end approach. This approach enabled different forwarding techniques, such as using path diversity—i.e. feeding a QKD service with several key streams—or combining keys from different vendors in each hop to provide a vendor-independent service.
Also, using the same approach, other quantum-enabled services have been trialed, such as quantum oblivious key distribution [58, 8] of quantum digital signature [59]. Of interest for this work is the first case, where some CV-QKD in the co-located domain were configured to deliver the raw key, without post-processing. Other DV-QKD still delivered symmetric distilled key. Using these two resources, key material suitable for quantum oblivious transfer could be obtained. This is an example of the SDN-aware interoperability needed to perform the end-to-end service. The SD-QKD Stack was modified to provide this service using the same software components and, as far as possible, the same standard interfaces. When not, the syntax of the SDN models was overloaded in order to contain the new configurations. Also, the configuration of operating parameters, such as wavelength, was trialed in Huawei CV-QKD systems using the QuAM/QuAI novel technology [60], so that the switching capabilities in the co-located domain could be configured using the same SD-QKD Stack.
VI Conclusion and future work
The MadQCI testbed not only addressed the challenge of installing 26 QKD multi-vendor modules in coexistence with classical equipment and communications, at 9 sites and 130 km of optical fibres. It has also addressed some high-level issues needed for real-world quantum networking such as the joint operation and management of classical and quantum resources; the integration in production sites of network operators and the fulfilment of their needs; the standardisation; or the quality assurance of the quantum-enhanced services provided.
In this way, it has been possible to present its results as a blueprint for large-scale quantum network deployments. The fundamentals of optical networking and quantum communications to support this goal have also been discussed. It has been important to characterise the networks in production, taking into account their complexity beyond the most basic components.
Indeed, MadQCI was deployed in a multi-domain scenario, with the support of REDIMadrid and Telefónica Innovación Digital. A multi-vendor quantum network—ID Quantique, Huawei, Toshiba, R&S and ADVA manufacturers—was installed with coexistence in production premises and real traffic examples. Several quantum and classical coexistence approaches were tested, ranging from more straightforward approaches to the testing of a switched QKD network. In all cases, the premises, infrastructure elements, classical equipment and information systems were industrial and standard solutions, suitable for deployment in production sites. Firmware updates were needed to take the QKD-generated keys into the encryptors. The performance of two of these approaches has been presented, demonstrating their feasibility for larger deployments.
The quantum connectivity achieved on each link has allowed to show how the SDN operation and management approach can be used to extend the performance to end-to-end quantum-enhanced services. The SD-QKD Stack enabled the testing of a large number of use cases. This software implemented several ETSI-GS-QKD standards, ensuring that it is a valid solution for deployments with a more ambitious scope. It is based on a centralised operation and management of resources through a logically centralized controller and a set of agents at each node, as well a set of local KMS that performs both the key management and networking functionalities. Thus, not only encryption applications have been supported, but also other less tested primitives, such as oblivious transfer.
MadQCI, is committed to be the reference instance in the deployment of other QCI and the pan-European quantum communication infrastructure EuroQCI. In fact, this network is now starting a new iteration thanks to investments related to the Quantum Flagship and Digital Europe Programme, as well as the MADQuantum-CM project [61], part of the Recovery and Resilience Facility funds managed by Spain. Design and procurement is underway and details are being finalised for collaboration with various partners to promote higher value-added use cases, including those involving the satellite segment.
Appendix A Schemes
This appendix gathers the most relevant nodal schemes. These reflect the application of the approach described in the text, as well as other details specific to each node. The nodes not depicted do not contain relevant details and their optics resemble the optics used to illuminate the fibre linking them. The legend used in figure 10 is the same for figures 11, 12, 13 and 14.
Figure 10: Layout in Norte node.
Figure 11: Layout in Concepción node.
Figure 12: Layout in Distrito node.
Figure 13: Layout in Quijote node.
Figure 14: Layout in Quevedo node.
Acknowledgments
The authors would like to thank projects MADQuantum-CM, funded by Comunidad de Madrid (Programa de Acciones Complementarias) and by the Recovery, Transformation and Resilience Plan—Funded by the European Union—(NextGenerationEU, PRTR-C17.I1); the EU Horizon Europe project ”Quantum Security Networks Partnership” (QSNP), grant 101114043; and EuroQCI-Spain, DEP grant 101091638.